Fiona and Hoggy gain a kilo each every week. Fiona is currently 70kg, while Hoggy is 10kg. How many weeks does it take for Fiona to be TWICE as heavy as Hoggy?
NOTE: I've asked nearly all my teachers in my school and none of them have gotten the answer without using trial and error. I want to know how to get the answer by using algebra. Thank you :)

Answer by Alan3354(69443)

Fiona and Hoggy gain a kilo each every week. Fiona is currently 70kg, while Hoggy is 10kg. How many weeks does it take for Fiona to be TWICE as heavy as Hoggy?
-------------
The difference remains 60 kg, so it's when Hoggy weighs 60 kg.
----
w = # of weeks
F = 70 + w
H = 10 + w
---
F = 2H
70 + w = 2(10+w)
70 + w = 20 + 2w
w = 50 weeks